# Context-Alignment: Activating and Enhancing LLM Capabilities in Time Series

### v2

@@ 1 Introduction
  • We emphasize that effectively leveraging LLMs for TS tasks requires first activating their capabilities and then enhancing them. Besides, we pinpoint that token-level alignment fails to fully activate pre-trained LLMs due to their neglect of LLMs’ inherent strengths, which primarily stem from a deep understanding of logic and structure, rather than superficial token processing.
  • We are the first to propose Context-Alignment paradigm, which aims to construct a context-level alignment between TS and language, thereby activating LLMs’ potential capabilities in TS tasks.
  • We develop a Dual-Scale Context-Alignment GNNs framework, which achieves structural and logical alignment through dual-scale nodes and directed edges, thus realizing Context-Alignment. Furthermore, by integrating the few-shot prompting technique, we introduce (FSCA), which enhances LLMs’ performance in TS tasks.
  • Our experiments across multiple datasets and various TS tasks demonstrate that our method surpasses existing techniques, especially in few-shot and zero-shot forecasting tasks. Ablation studies further emphasize the importance of Context-Alignment.
@@ 3.1 Preliminaries and Token Embedding
Time-Series Classification. Given TS data ${\bm{X}}\in\mathbb{R}^{D\times T}$, the task is to build a classification model $\mathcal{C}$ with parameters ${\bm{\Theta}}_{\mathcal{C}}$ and prompt ${\bm{P}}_{\mathcal{C}}$ to assign a class label $\hat{y}\in\{1,2,...,C\}$, where $C$ is the number of classes. The classification is defined as $\hat{y}=\mathcal{C}({\bm{X}};{\bm{\Theta}}_{\mathcal{C}},{\bm{P}}_{\mathcal{C}})$.
@@ 3.2 Demo: Vanilla Context-Alignment (VCA)
where ${\bm{\Gamma}}_{c\to f}\in\mathbb{R}^{2\times(n+m)}$ is a $0-1$ assignment matrix, ${\bm{\Gamma}}_{ij}=1$ means that the $j$-th node in $G_{F}$ be aggregated into the $i$-th node in $G_{C}$. ${\bm{W}}_{c\to f}\in\mathbb{R}^{M\times M}$ is a learnable weight matrix. $\hat{\bm{N}}_{F}$ can be updates as $\hat{\bm{N}}_{F}\leftarrow\hat{\bm{N}}_{F}+\Delta{\bm{N}}$ after aggregating the information from $\hat{\bm{N}}_{C}$.
@@ 3.3 Few-Shot Prompting Based Context-Alignment (FSCA)
In NLP, “few-shot prompting" refers to a small set of instances provided to the model to demonstrate a task, enabling the model to perform similar tasks effectively (Brown, 2020). Based on this, we divide the TS embeddings $\{\bm{e}_{i}\}_{i=1}^{n}$ into $N$ parts while preserving their original order. Since the subsequent TS part can be used as ground truth for predictions based on the preceding TS parts, we can construct $N-1$ prediction demonstration examples using $N$ parts of $\{\bm{e}_{i}\}_{i=1}^{n}$ and language prompt embeddings $\{{\bm{z}}_{1},{\bm{z}}_{2},\dots,{\bm{z}}_{m}\}$ same as demo in Sec. 3.2. The $j$-th part of $\{\bm{e}_{i}\}_{i=1}^{n}$ is denoted as $\{{\bm{e}}_{j,1},\dots,{\bm{e}}_{j,l_{j}}\}$. We arrange the TS-language embeddings in the format as:
@@ 4 Experiments
Baselines.  Building on the groundwork of Zhou et al. (2023) and Jin et al. (2024), and mindful of page constraints, we selected a representative array of high-performing baselines for extensive evaluation. These encompass Transformer-based models such as iTransformer (Liu et al., 2023a), FEDformer (Zhou et al., 2022), Non-stationary Transformer (Liu et al., 2022), ETSformer (Woo et al., 2022), PatchTST (Nie et al., 2022), alongside notable non-Transformer methods like TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023). We also included advanced LLM-based models—GPT4TS (Zhou et al., 2023), Time-LLM (Jin et al., 2024), and S2IP-LLM (Pan et al., 2024)—all utilizing GPT-2 as the standard LLM backbone to ensure model consistency. To ensure a fair comparison, we adhere to the experimental framework outlined in  Zhou et al. (2023) and  Wu et al. (2022). Detailed evaluations are expanded upon in subsequent sections.
@@ 4.1 Demo: Vanilla Context-Alignment
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[350]Q[110]Q[110]Q[110]Q[110], row1 = c, row2 = c, cell11 = r=2, cell12 = c=4, cell3,4,5,62,3,4,5 = c, hline1 = 1-60.15em, hline2 = 2-60.05em, hline3,5 = 1-60.05em, hline7 = 1-60.15em, vline2 = 1abovepos = -1.2, belowpos = -1.2, vline2,3,4,5 = 2abovepos = -1.2, belowpos = -1.2, vline2,3,4,5 = 3,5abovepos = -1.2,, vline2,3,4,5= 4,6belowpos = -1.2,, Method/Variant & Long-term Forecasting
@@ 4.2 Long-term Forecasting
{tblr}
width = 1colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], cells = c, cell11,3,5,7,9,11,13,15,17,19,21 = c=2, cell1,2,3,4,5,6,7,8,9,101 = c=2, hline1 = 1-240.15em, hline2,3 = 1-240.05em, hline11 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-9, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 10belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@ 4.3 Short-term Forecasting
{tblr}
width = 1colspec = Q[0.2em,c]Q[2em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c], colsep = 0.1pt, rowsep = 1.4pt, cell11 = c=20.09, cell21 = r=3, hline1 = 1-140.15em, hline2 = 1-140.05em, hline5 = 1-140.15em, vline3,4,5,6,7,8,9,10,11,12,13,14 = 3, vline3,4,5,6,7,8,9,10,11,12,13,14 = 2abovepos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 4belowpos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 1abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST N-HiTS N-BEATS TimesNet FEDformer Stationary
Average SMAPE 11.828 12.021 12.494 12.690 12.142 13.639 12.059 12.035 12.250 12.880 13.160 12.780
@@ 4.4 Few-shot Forecasting
{tblr}
width = 1colspec = Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], colsep = 2.2pt, rowsep = 1.4pt, cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell2,3,4,5,6,71 = c=2, hline1,8 = 1-240.15em, hline2,3,7 = 1-240.05em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-5, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 6belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2,7abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@
Setups.  The expressive potential of LLMs often results in a robust performance in few-shot scenarios (Brown, 2020; Liu et al., 2023b). While current LLM-based methods outperform earlier models like DLinear, PatchTST, and TimesNet, they do not fully exploit LLMs’ deep TS comprehension, thus not maximizing their few-shot capabilities. To explore whether Context-Alignment can boost the few-shot efficacy of LLMs, we perform experiments on the ETT datasets, following the protocol established by Jin et al. (2024). We evaluate the performance of FSCA using only 5% of training data, with results from 10% data detailed in Appendix C.3.
@@ 4.5 Zero-shot Forecasting
{tblr}
width = 1colsep = 2.2pt, rowsep = 1.4pt, cells = c, colspec = Q[150]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50]Q[50], cell12,4,6,8,10,12,14,16 = c=2, hline1 = 1-170.15em, hline2,3,4,5,6,7,8,9,10,11 = 1-170.05em, hline12 = 1-170.15em, vline2,4,6,8,10,12,14,16 = 1,2,3,4,5,6,7,8,9,10,11abovepos = -1.2, belowpos = -1.2,
Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet
@@ 4.7 Ablation Study
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[40]Q[40]Q[45]Q[45]Q[45]Q[45], row1,2,3 = c, cell11 = c=2,r=2, cell13,5 = c=2, cell23,4,5 = c, cell31,3,5 = c=2, cell4,5,6,7,8,9,10,11,12,13,14,15,16,173,4,5,6 = c, cell4,5,6,7,8,9,10,11,12,13,14,15,16,171 = c=2, hline1 = 1-60.15em, hline2 = 3-60.05em, hline3,4,6,7,13 = 1-60.05em, hline18 = 1-60.15em, vline3,5 = 1,2,3,6abovepos = -1.2, belowpos = -1.2, vline3,5 = 4,7,13abovepos = -1.2,, vline3,5 = 5,12,17belowpos = -1.2,, vline3,5 = 8-11,14-16, Variant & Long-term Forecasting Classification
@@ A.2 DATASET DETAILS
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[227]Q[169]Q[250]Q[250], cells = c, hline1,8 = 1-40.15em, hline2 = 1-40.05em, vline2,3,4 = 1abovepos = -1.2, belowpos = -1.2, vline2,3,4 = 2abovepos = -1.2, vline2,3,4 = 7belowpos = -1.2, vline2,3,4 = 3-6, Dataset & Length Dimension Frequency
@@
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[227]Q[127]Q[142]Q[188]Q[233], cells = c, cells = c, hline1,8 = 1-50.15em, hline2 = 1-50.05em, vline2,3,4,5 = 1abovepos = -1.2, belowpos = -1.2, vline2,3,4,5 = 2abovepos = -1.2, vline2,3,4,5 = 7belowpos = -1.2, vline2,3,4,5 = 3-6, Dataset & Length Horizon Frequency Information
@@
{tblr}
colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[277]Q[156]Q[144]Q[152]Q[94]Q[108], cells = c, hline1,12 = 1-60.15em, hline2 = 1-60.05em, vline2,3,4,5,6 = 1abovepos = -1.2, belowpos = -1.2, vline2,3,4,5,6 = 2abovepos = -1.2, vline2,3,4,5,6 = 11belowpos = -1.2, vline2,3,4,5,6 = 3-10, Dataset & Train Cases Test Cases Dimensions Length Classes
@@ A.3 BASELINE DETAILS
iTransformer [Liu et al., 2023a]: iTransformer repurposes the Transformer architecture by applying attention and feed-forward networks on inverted dimensions to improve multivariate time series forecasting, achieving state-of-the-art performance on real-world datasets.
@@ A.4 Evaluation Metrics
MASE $=\frac{1}{H}\sum_{h=1}^{H}\frac{|Y_{h}-\hat{Y}_{h}|}{\frac{1}{H-s}\sum_{j=s+1}^{H}|Y_{j}-Y_{j-s}|},$ OWA $=\frac{1}{2}\left[\frac{\text{SMAPE}}{\text{SMAPE}_{\text{Na\"{\i}ve2}}}+\frac{\text{MASE}}{\text{MASE}_{\text{Na\"{\i}ve2}}}\right],$
@@ C.1 Long-term Forecasting full results
{tblr}
width = 1colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell21 = c=2, cell3,8,13,18,23,28,33,381 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18,23,28,33,38 = 1-240.05em, hline43 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-6,9-11,14-16,19-21,24-26,29-31,34-36,39-41, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3,8,13,18,23,28,33,38abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 7,12,17,22,27,32,37,42belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@ C.2 Short-term Forecasting full results
{tblr}
width = 1colspec = Q[0.2em,c]Q[2em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c], colsep = 0.1pt, rowsep = 1.4pt, cells = c, cell11 = c=20.09, cell2,5,8,11,141 = r=3, hline1 = 1-140.15em, hline2,5,8,11,14 = 1-140.05em, hline17 = 1-140.15em, vline3,4,5,6,7,8,9,10,11,12,13,14 = 3,6,9,12,15, vline3,4,5,6,7,8,9,10,11,12,13,14 = 2,5,8,11,14abovepos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 4,7,10,13,16belowpos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 1abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST N-HiTS N-BEATS TimesNet FEDformer Stationary
Year. SMAPE 13.288 13.413 13.750 15.110 13.652 16.965 13.477 13.422 13.487 15.378 14.021 14.727
@@ C.3 Few-shot Forecasting full results
{tblr}
colspec = Q[1.5em,c]Q[2.5em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], colsep = 2.2pt, rowsep = 1.4pt, cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell21 = c=2, cell3,8,13,181 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18 = 1-240.05em, hline23 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-6,9-11,14-16,19-21, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3,8,13,18abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 7,12,17,22belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@
{tblr}
colspec = Q[1.5em,c]Q[2.5em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], colsep = 2.2pt, rowsep = 1.4pt, cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell21 = c=2, cell3,8,13,181 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18 = 1-240.05em, hline23 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-6,9-11,14-16,19-21, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3,8,13,18abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 7,12,17,22belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer

### v3

@@ 1 Introduction
$\bullet$ We emphasize that effectively leveraging LLMs for TS tasks requires first activating their capabilities and then enhancing them. Besides, we pinpoint that token-level alignment fails to fully activate pre-trained LLMs due to their neglect of LLMs’ inherent strengths, which primarily stem from a deep understanding of logic and structure, rather than superficial token processing.
$\bullet$ We are the first to propose Context-Alignment paradigm, which aims to construct a context-level alignment between TS and language, thereby activating LLMs’ potential capabilities in TS tasks.
$\bullet$ We develop a Dual-Scale Context-Alignment GNNs framework, which achieves structural and logical alignment through dual-scale nodes and directed edges, thus realizing Context-Alignment. Furthermore, by integrating the few-shot prompting technique, we introduce (FSCA), which enhances LLMs’ performance in TS tasks.
$\bullet$ Our experiments across multiple datasets and various TS tasks demonstrate that our method surpasses existing techniques, especially in few-shot and zero-shot forecasting tasks. Ablation studies further emphasize the importance of Context-Alignment.
@@ 3.1 Preliminaries and Token Embedding
[…]
@@ 3.2 Demo: Vanilla Context-Alignment (VCA)
where ${\bm{\Gamma}}_{c\to f}\in\mathbb{R}^{2\times(n+m)}$ is a 010-1 assignment matrix, ${\bm{\Gamma}}_{ij}=1$ means that the $j$-th node in $G_{F}$ be aggregated into the $i$-th node in $G_{C}$. ${\bm{W}}_{c\to f}\in\mathbb{R}^{M\times M}$ is a learnable weight matrix. $\hat{\bm{N}}_{F}$ can be updates as $\hat{\bm{N}}_{F}\leftarrow\hat{\bm{N}}_{F}+\Delta{\bm{N}}$ after aggregating the information from $\hat{\bm{N}}_{C}$.
@@ 3.3 Few-Shot Prompting Based Context-Alignment (FSCA)
In NLP, “few-shot prompting" refers to a small set of instances provided to the model to demonstrate a task, enabling the model to perform similar tasks effectively (Brown, 2020). Based on this, we divide the TS embeddings $\{\bm{e}_{i}\}_{i=1}^{n}$ into $N$ parts while preserving their original order. Since the subsequent TS part can be used as ground truth for predictions based on the preceding TS parts, we can construct N1N-1 prediction demonstration examples using $N$ parts of $\{\bm{e}_{i}\}_{i=1}^{n}$ and language prompt embeddings $\{{\bm{z}}_{1},{\bm{z}}_{2},\dots,{\bm{z}}_{m}\}$ same as demo in Sec. 3.2. The $j$-th part of $\{\bm{e}_{i}\}_{i=1}^{n}$ is denoted as $\{{\bm{e}}_{j,1},\dots,{\bm{e}}_{j,l_{j}}\}$. We arrange the TS-language embeddings in the format as:
@@ 4 Experiments
Baselines.  Building on the groundwork of Zhou et al. (2023) and Jin et al. (2024), and mindful of page constraints, we selected a representative array of high-performing baselines for extensive evaluation. These encompass Transformer-based models such as iTransformer (Liu et al., 2023b), FEDformer (Zhou et al., 2022), Non-stationary Transformer (Liu et al., 2022), ETSformer (Woo et al., 2022), PatchTST (Nie et al., 2022), alongside notable non-Transformer methods like TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023). We also included advanced LLM-based models—GPT4TS (Zhou et al., 2023), Time-LLM (Jin et al., 2024), and S2IP-LLM (Pan et al., 2024)—all utilizing GPT-2 as the standard LLM backbone to ensure model consistency. To ensure a fair comparison, we adhere to the experimental framework outlined in  Zhou et al. (2023) and  Wu et al. (2022). Detailed evaluations are expanded upon in subsequent sections.
@@ 4.1 Demo: Vanilla Context-Alignment
Method/Variant Long-term Forecasting
@@ 4.2 Long-term Forecasting
Methods FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@ 4.3 Short-term Forecasting
Methods FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST N-HiTS N-BEATS TimesNet FEDformer Stationary
Average
SMAPE 11.828 12.021 12.494 12.690 12.142 13.639 12.059 12.035 12.250 12.880 13.160 12.780
@@ 4.4 Few-shot Forecasting
Methods FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@
Setups.  The expressive potential of LLMs often results in a robust performance in few-shot scenarios (Brown, 2020; Liu et al., 2023a). While current LLM-based methods outperform earlier models like DLinear, PatchTST, and TimesNet, they do not fully exploit LLMs’ deep TS comprehension, thus not maximizing their few-shot capabilities. To explore whether Context-Alignment can boost the few-shot efficacy of LLMs, we perform experiments on the ETT datasets, following the protocol established by Jin et al. (2024). We evaluate the performance of FSCA using only 5% of training data, with results from 10% data detailed in Appendix C.3.
@@ 4.5 Zero-shot Forecasting
Methods FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet
@@ 4.7 Ablation Study
Variant Long-term Forecasting Classification
@@ A.2 DATASET DETAILS
Dataset Length Dimension Frequency
@@
Dataset Length Horizon Frequency Information
@@
Dataset Train Cases Test Cases Dimensions Length Classes
@@ A.3 BASELINE DETAILS
iTransformer [Liu et al., 2023b]: iTransformer repurposes the Transformer architecture by applying attention and feed-forward networks on inverted dimensions to improve multivariate time series forecasting, achieving state-of-the-art performance on real-world datasets.
@@ A.4 Evaluation Metrics
MASE $=\frac{1}{H}\sum_{h=1}^{H}\frac{|Y_{h}-\hat{Y}_{h}|}{\frac{1}{H-s}\sum_{j=s+1}^{H}|Y_{j}-Y_{j-s}|},$ OWA $=\frac{1}{2}\left[\frac{\text{SMAPE}}{\text{SMAPE}_{\text{Naïve2}}}+\frac{\text{MASE}}{\text{MASE}_{\text{Naïve2}}}\right],$
@@ C.1 Long-term Forecasting full results
Methods FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@ C.2 Short-term Forecasting full results
Methods FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST N-HiTS N-BEATS TimesNet FEDformer Stationary
Year.
SMAPE 13.288 13.413 13.750 15.110 13.652 16.965 13.477 13.422 13.487 15.378 14.021 14.727
@@ C.3 Few-shot Forecasting full results
Methods FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@
Methods FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer